# DSR-Bench: Evaluating the Structural Reasoning Abilities of LLMs via Data Structures

### v1

@@ Abstract
Large language models (LLMs) are increasingly deployed for real-world tasks that fundamentally involve data manipulation. A core requirement across these tasks is the ability to perform structural reasoning—that is, to understand and reason about data relationships. For example, customer requests require a temporal ordering, which can be represented by data structures such as queues. However, existing benchmarks primarily focus on high-level, application-driven evaluations without isolating this fundamental capability. To address this gap, we introduce DSR-Bench, a novel benchmark evaluating LLMs’ structural reasoning capabilities through data structures, which provide interpretable representations of data relationships. DSR-Bench includes 20 data structures, 35 operations, and 4,140 problem instances, organized hierarchically for fine-grained analysis of reasoning limitations. Our evaluation pipeline is fully automated and deterministic, eliminating subjective human or model-based judgments. Its synthetic nature also ensures scalability and minimizes data contamination risks. We benchmark nine state-of-the-art LLMs. Our analysis shows that instruction-tuned models struggle with basic multi-attribute and multi-hop reasoning. Furthermore, while reasoning-oriented models perform better, they remain fragile on complex and hybrid structures, with the best model achieving an average score of only 47% on the challenge subset. Crucially, models often perform poorly on multi-dimensional data and natural language task descriptions, highlighting a critical gap for real-world deployment.
@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive capabilities across practical tasks such as database management [62], coding [3, 58], and scheduling [20]. These tasks differ in domain, but they share a core requirement: the ability to understand and manipulate data according to relationships such as order, hierarchy, and connectivity—a capability we refer to as structural reasoning.
Structural reasoning is an essential, yet often implicit, requirement in problem solving. For instance, when using an LLM to plan a trip, the system may need to interpret a map of cities as a graph, construct an itinerary by finding a path through that graph, and prioritize visits using a queue. As LLMs are increasingly deployed in real-world applications, particularly safety-critical ones such as robotics [39], healthcare [38], and autonomous systems [14, 1], it becomes imperative to rigorously evaluate their structural reasoning capabilities to identify potential failure modes prior to deployment. Furthermore, LLMs are known to struggle with complex forms of reasoning, such as spatial reasoning [22, 55], underscoring the timely need for a systematic evaluation of their structural reasoning abilities.
However, existing benchmarks offer limited insight into structural reasoning. Most evaluations focus on high-level, domain-specific tasks such as mathematical problem solving [28, 25, 50], competitive programming [16], or STEM-related questions [15]. These tasks are typically multi-step and complex, making it difficult to isolate structural reasoning performance from other reasoning abilities. As a result, a critical gap remains: there are no targeted, fine-grained benchmarks for evaluating structural reasoning, which is a core skill underlying many downstream applications
To address this gap, we propose evaluating structural reasoning through the lens of data structures. Data structures offer a controlled and interpretable framework for modeling diverse data relationships: arrays for sequences, stacks and queues for temporal logic, trees for hierarchies, and graphs for complex networks. Since structural reasoning manifests as operations on these canonical structures, data structure tasks provide a principled way to probe LLM reasoning abilities at a granular level. Such a fine-grained evaluation can not only identify reliable applications of LLMs for safe deployment, but also pinpoint areas where reasoning breaks down and how it might be improved.
@@
Building on this insight, we introduce DSR-Bench, a benchmark designed to isolate and evaluate LLM structural reasoning via data structure manipulation tasks. DSR-Bench comprises 4,140 problem instances spanning 20 data structures (grouped into six relationship categories, see Figure 1) and 35 operations across three length types (short, medium, and long). We also introduce DSR-Bench-challenge, a subset of DSR-Bench consisting of data structures with particularly complex designs. Each task probes whether the model can construct, maintain, and reason about data relationships within an interpretable and automatically verifiable setting.
DSR-Bench offers several key advantages for evaluating structural reasoning:
Hierarchical organization. Tasks are organized by increasing structural complexity, enabling a fine-grained analysis of specific reasoning skills. Within each category, we design a range of tasks to isolate different sources of complexity, allowing structural reasoning to be broken down into progressively more challenging tasks. This approach precisely identifies the specific types of data relationships that pose difficulties for LLMs.
Deterministic evaluation. Each data structure task has a concise and well-defined final state as its ground-truth label, allowing for deterministic and unambiguous scoring. Unlike open-ended tasks, this design supports fully automated evaluation without the need for human or model-based judgment, resulting in a fairer and more objective evaluation pipeline [4, 10, 54].
We benchmark nine state-of-the-art LLMs, including both instruction-tuned and reasoning models. Our analysis reveals the following key insights:
1.
Instruction-tuned models struggle with multi-attribute and multi-hop reasoning. Their accuracy drops sharply on tasks involving multi-attribute data (such as hashmaps) and tasks requiring the maintenance of multi-hop properties (such as red-black trees). Prompting strategies that encourage intermediate steps can improve performance, but techniques like CoT tend to be most effective when the data structure is less standard.
2.
Reasoning models still struggle with complex structures and non-standard tasks. Despite outperforming instruction-tuned models, reasoning models perform poorly on structurally complex tasks, with the best scoring only 47% on DSR-Bench-challenge. Some also fail when task instructions conflict with what are likely learned assumptions (e.g., consistently predicting tree depth as the number of edges instead of nodes, despite explicit instructions). These errors suggest a reliance on entrenched priors, limiting the model’s ability to adapt to user-defined constraints.
3.
Model performance declines significantly on complex spatial data structures. Accuracy drops as data dimensionality increases and degrades further when data is non-uniform. These patterns suggest a reliance on memorization over robust generalization or genuine reasoning.
4.
LLMs perform worse when data structure tasks are presented in natural language. Translating formal task descriptions into natural language leads to a marked drop in performance, posing a barrier to the practical deployment of LLMs for real-world problem solving.
@@
We release all code at https://github.com/dransyhe/DSR-Bench and all data at https://huggingface.co/collections/vitercik-lab/dsr-bench-6826381f6297ff1499134163 to support full reproducibility. We welcome community engagement and collaboration.
@@ 2 DSR-Bench: The Data Structure Reasoning Benchmark
In DSR-Bench, we propose a taxonomy of data structures grouped into six categories based on the types of data relationships they encode: Linear, Temporal, Associative, Hierarchical, Network, and Hybrid. This taxonomy captures the diversity of relational patterns found in real-world data and enables systematic evaluation of LLM capabilities across distinct relationship types. Within each category, we include a spectrum of data structures that vary in complexity and operational constraints, allowing us to isolate specific reasoning challenges and precisely diagnose where models struggle.
2.1 Data structure tasks
@@
DSR-Bench includes the following data structures and tasks. A detailed description of each data structure is provided in Section A.2.
@@
Linear (Sequential): This category includes array and its operations: access, insert, delete, reverse, and search. Linear structures are fundamental as they introduce ordered relationships, enabling reasoning about position, sequence, and iteration, and they serve as a foundation for understanding more complex data abstractions.
@@
Temporal (Time-based ordering): Temporal structures include stack, queue, priority queue, and skip list, which operate under last-in-first-out, first-in-first-out, priority-based, or probabilistic rules. Their operations are compound actions of insertions and deletions. Temporal reasoning is essential in systems that require ordered execution over time, such as event queues and schedulers, and introduces an additional ordering constraint beyond simple linear sequencing.
@@
Associative (Key-value mapping): We evaluate hashmap, trie, and suffix tree through tasks focused on construction and compound insertion-deletion operations. These tasks test the ability to handle key-based insertions, retrievals, and pattern matching in nested or hashed structures, which are core skills for efficient lookup and structured access in systems such as databases.
@@
Hierarchical (Tree-like): This group includes binary search tree (BST), heap, red-black (RB) tree, and B+ tree, with tasks involving traversals and compound operations. These structures are common in file systems and database indexing. These tasks assess whether the model can maintain structural invariants, perform updates efficiently, and simulate recursive behavior.
@@
Network (Connectivity and group membership): This category includes Graph and Disjoint Set Union (DSU). Graph tasks focus on breadth- and depth-first traversal, testing a model’s ability to reason about connectivity and explore many-to-many relationships, such as user interactions in social networks. DSU tasks centre on union-find operations, which are fundamental for dynamically managing group membership and identifying connected components in partitioned data.
@@
Hybrid (Combined relationships): Real-world systems often rely on hybrid data structures that combine multiple relational principles. This category includes LRU cache, which integrates temporal and memory constraints; bloom filter, for probabilistic set membership; and Directed Acyclic Word Graph (DAWG), a trie-like hierarchical graph. Tasks involve compound insertions and deletions, testing whether models can reason over combined constraints and generalize beyond individual structural concepts.
2.2 Data generation
@@
We design a diverse set of operation tasks for each data structure, as detailed in Table 8. Tasks fall into three categories: construction, inspection (e.g., access, pre/in/post-order traversal), and manipulation (e.g., insert, delete). In addition to these atomic operations, which serve as fundamental building blocks and allow fine-grained diagnosis of specific failure modes, DSR-Bench also includes compound operations. These are composed of a sequence of manipulations (e.g., insert and delete) and are designed to test a model’s ability to maintain structural integrity across multiple steps.
All data is synthetically generated, which is efficient and less susceptible to training data contamination [56, 51]. Numerical inputs are sampled uniformly from the range 0 to 100, while string inputs are composed of lowercase English letters. For each task, ground-truth outputs are generated by implementing the corresponding data structure and executing the prescribed operations. The expected format of the final state is explicitly specified in the prompt. For simple structures like arrays, the final state is a sequence of elements. For more complex structures such as trees and graphs, the final state is represented using a canonical traversal, such as pre-order or breadth-first. The modular design of our benchmark also supports easy incorporation of new data structure types.
2.3 Suites
@@
We categorize each task into three levels based on input length: short (5–10), medium (11–20), and long (21–30), to evaluate length generalization. For atomic operations, length corresponds to the number of input elements, such as list items or tree nodes. For compound operations, tasks start from an empty structure, and length is defined by the number of sequential operations applied.
We also curated a DSR-Bench-challenge subset highlighting data structures with particularly complex designs, aimed at pushing the limits of state-of-the-art models. To further assess structural reasoning capabilities, we also introduce two extension suites: DSR-Bench-spatial, focusing on spatial reasoning on multi-dimensional inputs (details see Section 4.2), and DSR-Bench-natural, evaluating reasoning over natural language descriptions (details see Section 4.3).
@@ 2.4 Prompt design
For each task, we design a prompt template and populate it with synthetic data to produce individual problem instances. Each prompt follows a consistent format: (i) a concise description of the data structure; (ii) a detailed explanation of the operations to be performed, written to avoid ambiguity; (iii) the initial state of the data structure (e.g., an existing tree or list) and any additional inputs required to execute the task (e.g., new elements to insert or delete); and (iv) a specific question requesting the final computed outcome. Following recommended practices in prompt engineering [35], we also append the instruction “Answer the question in <token> tokens” to guide models toward producing concise outputs within a specified token budget. We also implement five different prompting methods for each task, as we detail in Section 3.2.
2.5 Structured output
A key advantage of our tasks is that model outputs can be deterministically verified using automated evaluation. This is possible because the final states of data structures can be represented as sequences of numbers or strings, enabling precise comparison. Unlike benchmarks involving proof writing or complex math problems, which often yield long and ambiguous outputs requiring subjective human or LLM-based judging [4, 10, 54], our evaluation framework is more objective and reproducible.
For automatic verification, models must produce outputs matching predefined JSON schemas enforced with Pydantic. For instance, a Binary Search Tree (BST) pre-order traversal returns a list[int]. We ensure that output formats are easy to parse: about 70% of the tasks yield a 1D integer array.
2.6 Scoring system
We evaluate each model by comparing its output to the ground-truth answer. Each correct output scores 1, otherwise 0. The overall performance is reported as the average accuracy across all categories and length levels. To ensure deterministic outputs and eliminate ambiguity during evaluation, every task admits a single, well-defined solution. Potential sources of nondeterminism, such as definitions of hash function and tie-breaking rules, are explicitly specified in the prompt and controlled during problem generation. Additionally, we run extensive tests across all tasks and manually inspect failure cases to verify that errors are not caused by ambiguity in the question statements or task specifications.
3 Experiment settings
In this section, we describe our experimental settings. Section 3.1 outlines the models evaluated, and Section 3.2 details the prompting methods used.
3.1 Models
We evaluate a broad set of LLMs across two categories: general-purpose instruction-tuned models and models designed for advanced reasoning. From each major provider, we select the latest publicly available flagship models. Each prompt is evaluated three times, and we report both the average and standard deviation, resulting in a total of 166,230 query evaluations.
Instruction-tuned LLMs are widely used in real-world applications due to their low latency, reduced cost, and scalability. Although not explicitly trained for algorithmic or mathematical reasoning, their ability to process structured data is critical in practical settings. We test GPT-4.1-2025-04-14 (OpenAI) [36], Llama3.3-70B (Meta) [13], Gemini-2.0-Flash-001 (Google DeepMind) [45] , Claude-3-5-Sonnet-20241022 (Anthropic) [2], and DeepSeek-V3 (DeepSeek) [7].
We also evaluate a suite of models optimized for advanced reasoning. These models are explicitly trained to handle complex, multi-step tasks by maintaining intermediate states, including: o4-mini-2025-04-16 (OpenAI) [34], Gemini-2.5-Pro-Preview-03-25 (Google DeepMind) [45], Claude-3-7-Sonnet-20250219 (Anthropic) [2], and DeepSeek-R1 (DeepSeek) [7].
3.2 Prompts
To examine how prompting strategies affect model performance on data structure tasks, we evaluate five configurations. Unlike reasoning models with internal mechanisms for multi-step inference (e.g., explicit reasoning tokens), instruction-tuned models are especially sensitive to prompt formulation. Our prompting configurations include: (i) Stepwise, which adds a “steps” field to the JSON output; (ii) 0-CoT, which appends “Let’s think step by step” without examples; (iii) CoT, which provides a single reasoning example with a final output; (iv) 3-shot, which includes three input-output examples; and (v) None, the default prompting setting with no added text. See Section A.3 for examples.

### v2

# Can LLMs Reason Structurally? An Evaluation via the Lens of Data Structures

@@ Abstract
As large language models (LLMs) take on increasingly complex tasks, understanding their algorithmic reasoning abilities has become essential. However, existing evaluations focus on distinct and isolated tasks. We propose a unified diagnostic lens: structural reasoning—understanding and manipulating relationships like order, hierarchy, and connectivity. We introduce DSR-Bench, the first benchmark to systematically evaluate LLM structural reasoning through canonical data structures, which serve as interpretable, algorithmically meaningful abstractions. DSR-Bench spans 20 data structures, 35 operations, and 4,140 synthetically generated problem instances with minimal contamination. The benchmark’s hierarchical design pinpoints specific failure modes, while its fully automated evaluation ensures objective and consistent assessment. Benchmarking ten state-of-the-art LLMs reveals critical limitations: the top-performing model scores only 0.498 out of 1 on challenging instances. Three additional evaluation suites reveal further weaknesses: models perform poorly on spatial data and natural language scenarios, and fail to reason over their own generated code. DSR-Bench offers a principled diagnostic tool for structural reasoning, helping expose reasoning bottlenecks and guide the development of more capable and reliable LLMs.
@@ 1 Introduction
As large language models (LLMs) tackle increasingly complex real-world challenges, strengthening their algorithmic reasoning abilities is essential for interpretability, safety, and efficiency (Eberle et al., 2025). Yet scaling alone is reaching its limits: models face data scarcity and diminishing returns (Villalobos et al., 2024). A complementary path is to study the algorithms LLMs learn and employ, which can reveal their internal representations and inspire algorithm-centric architectures that embed reasoning capabilities directly into model design (Bounsi et al., 2024; Eberle et al., 2025).
The call to prioritize algorithmic reasoning has intensified as the field shifts toward evaluating general reasoning without external tools. While LLMs can solve many tasks through code generation and execution, recent initiatives like Gemini-Deep-Think’s and OpenAI’s IMO competitions (Luong & Lockhart, 2025; Wei, 2025) explicitly prohibit coding and proof assistance. These efforts highlight a growing emphasis on end-to-end reasoning as a critical step toward artificial general intelligence.
Despite rapid progress, existing evaluations of algorithmic reasoning remain fragmented, each probing only a single facet, such as arithmetic operations (Zhou et al., 2022; 2024; Lee et al., 2024), simulation of textbook algorithms (Markeeva et al., 2024), or domain-specific tasks like graph problems (Wang et al., 2023a; Fatemi et al., 2024). Meanwhile, interpretability research, such as mechanistic interpretability (Olah et al., 2020; Wang et al., 2023b), examines internal model representations rather than reasoning over algorithmic structures. What is missing is a unified framework that bridges these threads, offering interpretable diagnostics of algorithmic reasoning across diverse problem types.
To address this gap, we propose to evaluate algorithmic reasoning through the lens of data structures—fundamental abstractions that organize information and govern algorithmic operations. Data structures provide an ideal evaluation framework because they are both systematic (covering diverse relationship types) and interpretable (offering clear diagnostic insights). They span the spectrum of relationships underlying algorithmic thinking: arrays represent sequences, stacks and queues capture temporal ordering, trees encode hierarchies, and graphs represent complex networks. This diversity enables us to holistically assess whether models can construct, maintain, and manipulate different types of relationships—the core capability we term structural reasoning.
Structural reasoning underlies many real-world applications. Trip planning requires interpreting maps as graphs and managing priorities with queues, while supply chain optimization depends on hierarchical resource allocation and temporal sequencing. In safety-critical domains such as robotics (Sado et al., 2023) and healthcare (Sadeghi et al., 2024), failures in structural reasoning abilities can have serious consequences, making systematic evaluation crucial before deployment.
Contributions
@@
We introduce DSR-Bench, the first benchmark to systematically isolate and evaluate LLM structural reasoning through data structure manipulation tasks. DSR-Bench comprises 4,140 problem instances spanning 20 data structures (grouped into six relationship categories; see Figure˜1) and 35 operations across three difficulty levels (short, medium, and long) and five evaluation suites (main, challenge, spatial, natural, and code). Each task provides interpretable, automatically verifiable assessments of whether models can construct, maintain, and reason about relationships. The highlights and main takeaways of DSR-Bench are:
Hierarchical and diagnostic design. Simpler tasks (e.g., queue operations) serve as prerequisites for complex ones (e.g., graph traversal). This enables fine-grained localization of failure modes.
Deterministic, contamination-resistant evaluation. Synthetic data minimizes contamination risk. The evaluation is deterministic and fully automated, supporting objective and fair assessment.
Five evaluation suites covering diverse reasoning settings.
main: Canonical data structure tasks reveal core reasoning gaps—e.g., instruction-tuned models struggle with multi-attribute and multi-hop reasoning, while reasoning models fail to override memory for non-standard tasks with user-defined constraints. Simpler prompts consistently help, but CoT requires careful design.
challenge: Hybrid and compositional structures remain difficult, with the best model scoring only 0.498 out of 1, revealing reasoning limitations of frontier models.
spatial: On multi-dimensional data, the common format found in real-world applications, models struggle as dimensionality increases or when data follows non-uniform distributions.
natural: In realistic natural language scenarios, models struggle to extract structure and navigate ambiguity, exposing a gap to real-world deployment.
code: An auxiliary probe contrasting reasoning with and without tool use. Models rarely benefit from reasoning over their own generated code; external interpreters help on familiar tasks but consistently fail on non-standard or natural-language variants.
Comprehensive empirical analysis. We evaluate ten state-of-the-art LLMs and analyze prompting strategies, distribution shifts, and qualitative error patterns (e.g., implicit priors, instruction-following failures). These findings provide actionable insights for developing architectures and training methods that better capture algorithmic reasoning.
@@
To welcome community engagement and collaboration, we release all code at https://github.com/dransyhe/DSR-Bench and data at https://huggingface.co/collections/vitercik-lab/dsr-bench-6826381f6297ff1499134163 for full reproducibility.
2 Related work
Algorithmic reasoning with LLMs
Prior works target distinct but isolated facets of algorithmic reasoning, often limited in scope: arithmetic tasks and length generalization in transformers (Zhou et al., 2022; 2024; Lee et al., 2024), or specific graph problems like cycle detection and connectivity (Wang et al., 2023a; Fatemi et al., 2024). CLRS-Text (Markeeva et al., 2024) evaluates whether LLMs can simulate 30 classical algorithms (Veličković et al., 2022; Cormen et al., 2009). We take a holistic and foundational approach using data structures, which serve as interpretable building blocks of algorithms and support fine-grained diagnosis of specific reasoning failures to inform future research.
Reasoning benchmarks
Existing LLM reasoning benchmarks are predominantly high-level and domain-specific, targeting math (Cobbe et al., 2021; Liu et al., 2024a; b), STEM (Hendrycks et al., 2021), and logic puzzles (White et al., 2025; Giadikiaroglou et al., 2024). These often require complex responses with intertwined reasoning steps, relying on subjective human or LLM-based evaluation (Chiang et al., 2024; Feuer et al., 2024; Ye et al., 2024). We focus on structural reasoning, an implicit requirement underlying problem-solving across domains. By using data structures as clear abstractions of these relationships, we isolate structural reasoning from domain-specific complexities.
Coding benchmarks
Coding benchmarks evaluate how well LLMs write syntactically correct code or function as coding agents (Chen et al., 2021; Zheng et al., 2023; Jimenez et al., 2024; Jain et al., 2025; White et al., 2025; Aider-AI, 2025), typically requiring external interpreters for verification. While useful, these benchmarks conflate reasoning with tool execution and are limited to domains where coding applies. In contrast, we target general reasoning independent of external tools, reflecting the broader goal of assessing progress toward general intelligence. To bridge these perspectives, we include the code suite to probe whether code generation aids structural reasoning.
@@ 3 DSR-Bench: the Data Structure Reasoning Benchmark
In DSR-Bench, we propose a taxonomy of data structures grouped into six categories based on the types of data relationships they encode: Linear, Temporal, Associative, Hierarchical, Network, and Hybrid. This taxonomy captures the diversity of relational patterns found in real-world data, enabling systematic evaluation across distinct reasoning challenges.
3.1 Tasks
@@
DSR-Bench includes the following data structures and tasks, further detailed in Section˜A.2.
@@
Linear (Sequential): This category includes array and its operations: access, insert, delete, reverse, and search. Linear structures introduce ordered relationships, enabling reasoning about position, sequence, and iteration. They serve as a foundation for more complex data abstractions.
@@
Temporal (Time-based ordering): Temporal structures include stack, queue, priority queue, and skip list, which operate under last-in-first-out, first-in-first-out, priority-based, or probabilistic rules. Their operations are compound actions of insertions and deletions. Temporal reasoning is essential in systems that require ordered execution over time, such as event queues and schedulers.
@@
Associative (Key-value mapping): We evaluate hashmap, trie, and suffix tree through tasks focused on construction and compound insertion-deletion operations. Tasks test the ability to handle key-based insertions, retrievals, and pattern matching in nested or hashed structures. These skills are core for efficient lookup and structured access in systems such as databases.
@@
Hierarchical (Tree-like): This group includes binary search tree (BST), heap, red-black (RB) tree, and B+ tree, with tasks involving traversals and compound operations. These structures are common in file systems and database indexing. Tasks assess whether the model can maintain structural invariants, perform updates efficiently, and simulate recursive behavior.
@@
Network (Connectivity and group membership): This category includes Graph and Disjoint Set Union (DSU) tasks. Graph tasks cover breadth-first and depth-first traversal, testing the ability to reason about connectivity and explore many-to-many relationships, as commonly found in social networks. DSU tasks evaluate union–find operations, which are essential for network connectivity analysis and clustering algorithms.
@@
Hybrid (Combined relationships): Real-world systems often rely on hybrid data structures that combine multiple forms of structural reasoning. This category includes LRU cache, which integrates temporal and memory management; bloom filter, for probabilistic set membership; and Directed Acyclic Word Graph (DAWG), a trie-like hierarchical graph. Tasks test whether models can compose different principles and generalize beyond individual structures.
Hierarchical organization
Tasks are organized by increasing complexity for fine-grained diagnosis of reasoning failures. For example, success on list but failure on hashmap isolates the issue to multi-attribute access rather than linear sequencing. Similarly, since tree traversal depends on queue or stack operations, testing these primitives becomes a prerequisite.
Operations
@@
We design a diverse set of operation tasks for each data structure (full list in Table˜10), spanning three categories: construction, inspection (e.g., access, traversal), and manipulation (e.g., insert, delete). Beyond these atomic operations, DSR-Bench includes compound operations: sequences of manipulations (e.g., insert, insert, delete, …) designed to test whether models can maintain structural integrity across multiple steps.
Difficulty levels
@@
We assign tasks to three difficulty levels based on input length: short (5–10), medium (11–20), and long (21–30), to assess length generalization. For atomic operations, length refers to the number of input elements (e.g., list items, tree nodes). For compound operations, tasks begin with an empty structure, and length is defined by the number of sequential operations.
Evaluation suites
Beyond the main suite, we curated a challenge subset targeting particularly complex structures to stress-test advanced reasoning capabilities. To more comprehensively evaluate structural reasoning beyond canonical data structures, we introduce three additional suites: spatial, which evaluates reasoning on multi-dimensional data; natural, which tests reasoning when tasks are embedded in realistic natural language scenarios; and code, which assesses whether models can leverage code generation for structural reasoning. Each suite examines how well structural reasoning capabilities transfer to scenarios critical for real-world deployment.
@@ 3.2 Prompt design
For each task, we design a prompt template and populate it with synthetic data to produce individual problem instances. Each prompt follows a consistent format: (i) a concise description of the data structure; (ii) a detailed explanation of the operations to be performed, written to avoid ambiguity; (iii) the initial state of the data structure (e.g., an existing tree or list) and any additional inputs required to execute the task (e.g., new elements to insert or delete); and (iv) a specific question requesting the final computed outcome. Following recommended practices in prompt engineering (OpenAI, 2025), we also append the instruction “Answer the question in <number> tokens” to guide models toward producing concise outputs within a specified token budget. We also implement five different prompting strategies for each task, as we detail in Section˜A.3.
3.3 Pipeline
Data generation
All data is synthetically generated to ensure efficiency, scalability, and minimal risk of training data contamination (Zhang et al., 2024a; Xu et al., 2024a). Each data structure and its operations are programmatically implemented to produce ground-truth outputs. Inputs are randomly sampled—numbers uniformly from 0–100 and strings from lowercase English letters. This modular approach enables accurate, reproducible evaluation and easy integration of new data structure types.
Automated evaluation
The evaluation pipeline is fully objective and reproducible through automated and deterministic verification. This avoids subjective judging in proof-writing or math benchmarks (Chiang et al., 2024; Feuer et al., 2024; Ye et al., 2024). We integrate Structured Output using a predefined JSON Schema (natively supported by most models or via the Instructor library) and Pydantic, ensuring strict schema adherence. For example, a BST pre-order traversal must return a list[int]. Across 1,624 trials with nine models and six structures, we observed zero schema violations (Table˜33 in Appendix), confirming robustness of Structured Output.
Scoring system
We compare model outputs to ground-truth answers, scoring 1 for correct and 0 for incorrect responses. Overall performance is reported as average accuracy across categories and difficulty levels from three runs. To ensure determinism, each task has a single well-defined solution. Any potential ambiguity (hash functions, tie-breaking rules) is explicitly specified in the prompts. We include an ablation with Levenshtein distance as a partial scoring metric in Section˜A.10.